A used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized sedans was obtained. TheCalculated Numeric: A used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized sedaCalculated Numeric: A used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized sedaA used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized sedans was obtained. TheCalculated Numeric: A used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized sedaCalculated Numeric: A used car salesman wants to explain car price ($1,000s) using car age (years). A sample of midsized seda
Answer #1

Ho: B1 - 0 vs. Ha: B10

-----------------

strenght = Coefficient of regression

= sqrt(r square)

=0.9089

---------------------

Price depreciates by

= 1.2283 * 1000

= $1228.3

--------------------------

t-test statistic =    t = estimated slope / std error =   -1.2283   /   0.213   =   -5.767

-----------

y = 17.37 - 2*1.2283

=14.9134

------------

82.61% variability

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!